# A novel characterization of structures in smooth regression curves: from a viewpoint of persistent homology

### v3

@@ 7 Conclusion
In this article, we explored structures such as monotonicity, convexity, and modality in smooth regression curves via persistent homology of the super-level sets of a particular function. That particular function is the first derivative of the regression function. In general, persistent homology can be applied to identify structures in level sets of functions in terms of connected components and holes. We exploit the fact that derivatives of a function carry important geometric information about the function; therefore, one can identify useful structures in a function by investigating the persistent homology of its derivatives. The main problem that impedes one from using persistent homology to explore the geometric properties of a function via its derivative is the consistency of the estimates of persistent homology of its derivatives. The problem of consistency for the estimates of persistent homology arises due to the non-availability of an estimator of the derivative of the function that converges uniformly. Therefore, the usual procedure for estimating the persistent homology is ruled out in this case.
@@
However, [1] proposed a different procedure based on the kernel estimation of statistical functions (density and regression functions), applicable in the present context. We apply this procedure to estimate the persistent homology of the first derivative of regression functions and establish its consistency. This procedure is useful in the present context since its consistency does not require uniform convergence of the estimator of the underlying function. Thus, this procedure paves the way for exploring structures via the persistent homology of the first derivative of the regression functions. For future considerations, it would be tempting to extend this procedure for higher-order derivatives to explore some stimulating structures in regression curves. We would also like to explore the possibility of extending this procedure for the gradient of the regression function when there is more than one regressor. In this context, we would like to discuss the possibility of extending the procedure when there is more than one regressor and the difficulties involved therein. Note that when there is more than one regressor, the regression function is $m:\mathbb{R}^{d}\xrightarrow{}\mathbb{R},d\geq 2$. In this case, the first derivative, that is the gradient of the function, will be a function $m_{1}:\mathbb{R}^{d}\xrightarrow{}\mathbb{R}^{d},d\geq 2$. The usual persistent homology, also called one-parameter persistent homology, tracks the evolution of homological features in nested sequences of topological spaces indexed by a parameter $t\in\mathcal{I}\subset\mathbb{R}$. For example, filtrations built on point cloud usually have a radius $r\geq 0$ of Euclidean balls, or in super level-set filtrations, usually, it is the height of a real-valued function that varies over $\mathbb{R}$. Therefore, the usual persistent homology can be used only for real-valued functions. In contrast, here we aim to compute the persistent homology of a vector-valued function that does not have the usual notion of height. Thus, to compute persistent homology for the first derivative of a function when the domain of the function is $\mathbb{R}^{d},d\geq 2$, we have to resort to the so-called, multiparameter persistent homology. The multiparameter persistent homology is meant to track the evolution of topological features for filtrations of the form $\{\mathcal{X}_{\boldsymbol{t}};\boldsymbol{t}\in\mathbb{R}^{d},\},d\geq 2$. However, it does not have a straightforward extension of one-parameter persistent homology because it does not have a nice numerical invariant such as barcode. Thus, though it provides some information about the underlying topological space, it is not a compact numerical invariant, making it restrictive for topological inference. For these reasons, it is not so straightforward to make topological inferences about the derivative of the regression function when there is more than one regressor.
@@ 8.2 Definitions
First, we state and prove the following two lemmas that will be used to prove the Theorem 3.1. Then we provide proof of the Theorem 3.1. The Lemma 8.1 and Lemma 8.2 can be interpreted as follows. First, recall the set $\mathcal{X}^{t}_{n}\triangleq\{X_{i}:\hat{f}_{n}(X_{i})\geq t,i=1,\ldots,n\}$ in step 2 of the described procedure. Note that even small errors in the estimator $\hat{f}_{n}$ may result in a false positive or false negative error in the assignments of data points to $\mathcal{X}^{t}_{n}$ with respect to the super-level set that corresponds to the true function $m_{1}$. The true super-level set is denoted as $D_{t}$ here. A tiny error in $\mathcal{X}^{t}_{n}$ may introduce large errors in the estimated persistent homology of $D_{t}$. Therefore we need to find the probability of false positive and false negative assignments of data points to the set $\mathcal{X}^{t}_{n}$ with respect to $D_{t}$. We find an upper bound of false positive and false negative errors by inflating and deflating the set $D_{t}$ by a radius $h$. In other words, we define the two sets $D_{t}^{+}(h)$ and $D_{t}^{-}(h)$ as follows. First, we define the tube of radius of $h$ around the boundary of $D_{t}$, That is
$\partial D_{t}(h)\triangleq\bigcup_{x\in\partial D_{t}}\mathbb{B}_{h}(x),\partial D_{t}\text{ is boundary of }D_{t}.$
@@
$D_{t}^{+}(h)\triangleq D_{t}\cup\partial D_{t}(h)\text{ and }D_{t}^{-}(h)\triangleq D_{t}\setminus\partial D_{t}(h),$
where $D_{t}^{+}(h)$ and $D_{t}^{-}(h)$ corresponds to inflating and deflating the set $D_{t}$ by a radius $h$, respectively. The Lemma 8.1 computes upper bounds for false positive and false negative errors of assignments of data points to the set $\mathcal{X}_{n}^{t}$. Then using the Lemma 8.1, Lemma 8.2 shows that the set $\hat{D}_{t}(n,h)$ (See, Equation 3.4) is sandwiched between two non-random approximations of $D_{t}$ with high probability.
Lemma 8.1
Under the assumptions (A.1)-(A.5) and (B.1)-(B.5), for every $t>0$ and $\epsilon\in(0,t)$, if $h\equiv h(n)\xrightarrow{}0\text{ as }n\xrightarrow{}\infty$ such that $nh^{2}\xrightarrow{}\infty$, then there exists a constant $C_{\epsilon}$ such that for large enough $n$ we have
$\mathbb{P}\left(\exists X_{i}\in\mathcal{X}_{n}^{t}|X_{i}\notin D_{t-\epsilon}^{+}(h)\right)\leq ne^{-C_{\epsilon}nh^{2}}$ (8.1)
@@
$\mathbb{P}\left(\exists X_{i}\notin\mathcal{X}_{n}^{t}|X_{i}\in D_{t+\epsilon}^{-}(h)\right)\leq ne^{-C_{\epsilon}nh^{2}},$ (8.2)
@@
Now, we shall use the strong consistency of the Nadaraya-Watson estimator of $m$, denoted as $\hat{m}_{n}$. We have
$\hat{m}_{1,n}(x)$ $=\frac{\sum\limits_{i=1}^{n}K_{1,h}\left(x-X_{i}\right)Y_{i}}{\sum\limits_{i=1}^{n}K_{h}(x-X_{i})}-\frac{\sum\limits_{i=1}^{n}K_{1,h}\left(x-X_{i}\right)\hat{m}_{n}(x)}{\sum\limits_{i=1}^{n}K_{h}(x-X_{i})}$
$\asymp\frac{\sum\limits_{i=1}^{n}K_{1,h}(x-X_{i})\left(Y_{i}-m(x)\right)}{\sum\limits_{i=1}^{n}K_{h}(x-X_{i})}\text{ almost surely },$ (E.3)
where E.3 follows from the fact that $\hat{m}_{n}(x)\asymp m(x)\text{ almost surely for all }x\in\mathbb{R}$.
@@
$\asymp\mathbb{P}\left[\frac{\sum\limits_{i=1}^{n}K_{1,h}\left(X_{1}-X_{i}\right)\left(Y_{i}-m(X_{1})\right)}{\sum\limits_{i=1}^{n}K_{h}\left(X_{1}-X_{i}\right)}\geq t|X_{1}=x,Y_{1}=y\right]$ (Using E.3)
$=\mathbb{P}\left[\sum\limits_{i=2}^{n}\left\{K_{1,h}\left(x-X_{i}\right)\left(Y_{i}-m(x)\right)-tK_{h}\left(x-X_{i}\right)\right\}\geq t\right]$
@@
$\lessapprox\mathbb{P}\left[\sum\limits_{i=2}^{n}\left(Z_{i}-\mathbb{E}\left(Z_{i}\right)\right)\geq t+\epsilon\delta p_{min}2(n-1)h\right],$ (E.4 )
@@
$Z_{i}=K_{1,h}\left(x-X_{i}\right)\left(Y_{i}-m(x)\right)-(m_{1}(X_{i})+\epsilon)K_{h}\left(x-X_{i}\right).$ (8.3)
@@
Now, we would like to apply Hoeffding’s inequality (See Definition 8.1) for the random variables $Z_{i}-\mathbb{E}Z_{i},i=2,\ldots n$, therefore consider the following:
@@
    We have $|Z_{i}|\leq\tau\left(Y_{max}+M\right)+(M_{1}+\epsilon)$
Therefore, using 1 and 2 above, and applying Hoeffding’s inequality for $u=t+\epsilon\delta p_{min}2(n-1)h$, we have
$\mathbb{P}\left[\hat{m}_{1,n}(X_{1})\geq t|X_{1}=x,Y_{1}=y\right]$ $\leq\exp{\frac{-2\left(t+\epsilon\delta p_{min}2(n-1)h\right)^{2}}{4\left(n-1\right)\left[\tau\left(Y_{max}+M\right)+(M_{1}+\epsilon)\right]^{2}}}$
$=\exp{\frac{-n^{2}h^{2}\left(\frac{t}{nh}+\epsilon\delta p_{min}2\frac{\left(n-1\right)}{n}\right)^{2}}{2(n-1)\left[\tau\left(Y_{max}+M\right)+(M_{1}+\epsilon)\right]^{2}}}$
$\asymp\exp{\frac{-nh^{2}2\epsilon^{2}\delta^{2}p_{min}^{2}}{\left[\tau\left(Y_{max}+M\right)+(M_{1}+\epsilon)\right]^{2}}}$
$=\exp{(-nh^{2}C_{\epsilon})},$ (E.5)
where $C_{\epsilon}$ is a constant depending on $\epsilon$, defined as
$C_{\epsilon}=\displaystyle{\frac{2\epsilon^{2}\delta^{2}p_{min}^{2}}{\left[\tau\left(Y_{max}+M\right)+(M_{1}+\epsilon)\right]^{2}}}.$ (8.4)
Now, using the Equation E.5 in the Equation E.2 we have
[…]
$=ne^{-nh^{2}C_{\epsilon}}$
@@
$Z_{i}^{*}=\left(m_{1}(X_{i})-\epsilon\right)K_{h}(x-X_{i})-K_{1,h}(x-X_{i})\left(Y_{i}-m(x)\right).$
@@
$\mathbb{P}\left[\hat{m}_{1,n}(X_{1})\leq t|X_{1}=x,Y_{1}=y\right]\leq\mathbb{P}\left[\sum\limits_{i=2}^{n}\left(Z_{i}^{*}-\mathbb{E}\left(Z_{i}^{*}\right)\right)\geq\epsilon\delta p_{min}2(n-1)h-t\right].$
@@
This completes the proof of the Lemma 8.1. $\Box$
Lemma 8.2
Under the assumptions (A.1)-(A.5) and (B.1)-(B.5), for every $t>0$, and $\epsilon\in(0,t)$ if $h\equiv h(n)\xrightarrow{}0\text{ as }n\xrightarrow{}\infty\text{ such that }nh^{2}\xrightarrow{}\infty$, then for large enough $n$, we have
$\mathbb{P}(D^{-}_{t+\epsilon}(2h)\subset\hat{D}_{t}(n,h)\subset D^{+}_{t-\epsilon}(2h))\geq 1-3ne^{-C_{\epsilon}nh^{2}}.$ (8.5)
@@
Now, since $\hat{D}_{t}(n,h)\supset\mathcal{X}_{n}^{t}$, using the Lemma 8.1 we have
@@
$\leq\mathbb{P}\left(\mathcal{X}_{n}^{t}\cap\left(D^{+}_{t-\epsilon}(h)\right)^{c}\neq\varnothing\right)\leq ne^{-nh^{2}C_{\epsilon}}.$ (E.6)
@@
[…]
@@
$\mathbb{P}\left(\exists u\in\mathcal{I}:\mathbb{B}_{(1-\gamma)h}(u)\cap\mathcal{X}^{t}_{n}=\varnothing\right)\leq N_{\gamma}2\gamma he^{-n2(1-\gamma)h^{2}p_{min}}=N_{\gamma}2\gamma he^{-nh^{2}C_{\gamma}},$ (8.9)
where $C_{\gamma}=2(1-\gamma)p_{min}$. Now, note that the second term in Equation E.8 can be bounded by Lemma 8.1. Therefore, using the Equation 8.2, we have the following bound for the second term of Equation E.8,
$\mathbb{P}\left(D^{-}_{t+\epsilon}(h)\cap\mathcal{X}_{n}\not\subset\mathcal{X}^{t}_{n}\right)\leq ne^{-C_{\epsilon}nh^{2}}$ (8.10)
Now, using Equation 8.9 and Equation 8.10 in the Equation E.8, we have
$\mathbb{P}\left(E_{1}^{c}\right)\leq N_{\gamma}2\gamma he^{-nh^{2}C_{\gamma}}+ne^{-C_{\epsilon}nh^{2}}\leq 2ne^{-C_{\epsilon}nh^{2}},$ (8.11)
@@
Thus, using Equation 8.11 and Equation E.6 in Equation 8.6, we get the desired inequality as in Equation 8.5. This completes the proof of Lemma 8.2. $\Box$
@@
Therefore, it is enough to prove that $\delta_{B}\left(\widehat{PH}_{k}^{\epsilon}(m_{1}),PH_{k}^{\epsilon}(m_{1})\right)\leq 3\epsilon$ holds with high probability. Let E be the event that the following sequence of inclusions holds:
$D_{t_{0}+\epsilon}\hookrightarrow\hat{D}_{t_{0}}(n,h)\hookrightarrow D_{t_{1}+\epsilon}\hookrightarrow\hat{D}_{t_{1}}(n,h)\hookrightarrow D_{t_{2}+\epsilon}\hookrightarrow\cdots$
Now, applying Lemma 8.2 $N_{\epsilon}$ times for $\hat{D}_{t_{i}}(n,h)$, we have
$\mathbb{P}(E)\geq 1-3nN_{\epsilon}e^{-C_{\epsilon/2}nh^{2}}$
@@
$\mathds{P}\left(\delta_{B}\left(\widehat{PH}_{k}^{\epsilon}\left(m_{1}\right),PH_{k}\left(m_{1}\right)\right)\leq 5\epsilon\right)\geq 1-3nN_{\epsilon}e^{-C_{\epsilon/2}nh^{2}}.$

### v4

@@ 7 Conclusion
In this article, we explored structures such as monotonicity, convexity, and modality in smooth regression curves via persistent homology of the super-level sets of a particular function. That particular function is the first derivative of the regression function. In general, persistent homology can be applied to identify structures in level sets of functions in terms of connected components and holes. We exploit the fact that derivatives of a function carry important geometric information about the function, therefore one can identify useful structures in a function by investigating the persistent homology of its derivatives. The main problem that impedes one from using persistent homology to explore the geometric properties of a function via its derivative is the consistency of the estimates of persistent homology of its derivatives. The problem of consistency for the estimates of persistent homology arises due to the non-availability of an estimator of the derivative of the function which converges uniformly. Therefore, the usual procedure for estimating the persistent homology is ruled out in this case.
However, [1] proposed a different procedure based on the kernel estimation of statistical functions (density and regression functions), which is applicable in the present context. We apply this procedure to estimate the persistent homology of the first derivative of regression functions and establish its consistency. This procedure is useful in the present context since its consistency does not require uniform convergence of the estimator of the underlying function. Thus, this procedure paves the way for exploring structures via the persistent homology of the first derivative of the regression functions. For future considerations, it would be tempting to extend this procedure for higher-order derivatives to explore some stimulating structures in regression curves.
@@
We would also like to explore the possibility of extending this procedure for the gradient of the regression function when there is more than one regressor. In this context, we would like to discuss the possibility of extending the procedure when there is more than one regressor and the difficulties involved therein. Note that when there is more than one regressor, the regression function is $m:\mathbb{R}^{d}\xrightarrow{}\mathbb{R},d\geq 2$. In this case, the first derivative, that is the gradient of the function, will be a function $m_{1}:\mathbb{R}^{d}\xrightarrow{}\mathbb{R}^{d},d\geq 2$. The usual persistent homology, also called one-parameter persistent homology, tracks the evolution of homological features in nested sequences of topological spaces indexed by a parameter $t\in\mathcal{I}\subset\mathbb{R}$. For example, filtrations built on point cloud usually have a radius $r\geq 0$ of Euclidean balls, or in super level-set filtrations, usually, it is the height of a real-valued function that varies over $\mathbb{R}$. Therefore, the usual persistent homology can be used only for real-valued functions. In contrast, here we aim to compute the persistent homology of a vector-valued function that does not have the usual notion of height. Thus, to compute persistent homology for the first derivative of a function when the domain of the function is $\mathbb{R}^{d},d\geq 2$, we have to resort to the so-called, multiparameter persistent homology. The multiparameter persistent homology is meant to track the evolution of topological features for filtrations of the form $\{\mathcal{X}_{\boldsymbol{t}};\boldsymbol{t}\in\mathbb{R}^{d},\},d\geq 2$. However, it does not have a straightforward extension of one-parameter persistent homology because it does not have a nice numerical invariant such as barcode. Thus, though it provides some information about the underlying topological space, it is not a compact numerical invariant, making it restrictive for topological inference. For these reasons, it is not so straightforward to make topological inferences about the derivative of the regression function when there is more than one regressor.
@@ 8.2 Definitions
[…]
First, we state and prove the following two lemmas that will be used to prove the Theorem 3.1. Then we provide proof of the Theorem 3.1. The Lemma 1 and Lemma 2 can be interpreted as follows. First, recall the set $\mathcal{X}^{t}_{n}\triangleq\{X_{i}:\hat{f}_{n}(X_{i})\geq t,i=1,\ldots,n\}$ in step 2 of the described procedure. Note that even small errors in the estimator $\hat{f}_{n}$ may result in a false positive or false negative error in the assignments of data points to $\mathcal{X}^{t}_{n}$ with respect to the super-level set that corresponds to the true function $m_{1}$. The true super-level set is denoted as $D_{t}$ here. A tiny error in $\mathcal{X}^{t}_{n}$ may introduce large errors in the estimated persistent homology of $D_{t}$. Therefore we need to find the probability of false positive and false negative assignments of data points to the set $\mathcal{X}^{t}_{n}$ with respect to $D_{t}$. We find an upper bound of false positive and false negative errors by inflating and deflating the set $D_{t}$ by a radius $h$. In other words, we define the two sets $D_{t}^{+}(h)$ and $D_{t}^{-}(h)$ as follows. First, we define the tube of radius of $h$ around the boundary of $D_{t}$, That is
[…]
@@
[…]
where $D_{t}^{+}(h)$ and $D_{t}^{-}(h)$ corresponds to inflating and deflating the set $D_{t}$ by a radius $h$, respectively. The Lemma 1 computes upper bounds for false positive and false negative errors of assignments of data points to the set $\mathcal{X}_{n}^{t}$. Then using the Lemma 1, Lemma 2 shows that the set $\hat{D}_{t}(n,h)$ (See, Equation 3.4) is sandwiched between two non-random approximations of $D_{t}$ with high probability.
Lemma 1
Under the assumptions (A.1)-(A.5) and (B.1)-(B.5), for every $t>0$ and $\epsilon\in(0,t)$, if $h\equiv h(n)\xrightarrow{}0\text{ as }n\xrightarrow{}\infty$ such that $nh^{6}\xrightarrow{}\infty$, then there exists a constant $C_{\epsilon}$ such that for large enough $n$ we have
$\mathbb{P}\left(\exists X_{i}\in\mathcal{X}_{n}^{t}|X_{i}\notin D_{t-\epsilon}^{+}(h)\right)\leq ne^{-C_{\epsilon}nh^{6}}$ (8.1)
@@
$\mathbb{P}\left(\exists X_{i}\notin\mathcal{X}_{n}^{t}|X_{i}\in D_{t+\epsilon}^{-}(h)\right)\leq ne^{-C_{\epsilon}nh^{6}},$ (8.2)
@@
Now, consider the following:
$\hat{m}_{1,n}(x)$ $=\frac{\sum\limits_{i=1}^{n}K_{1,h}(x-X_{i})\left(Y_{i}-\hat{m}_{n}(x)\right)}{\sum\limits_{i=1}^{n}K_{h}(x-X_{i})}$
[…]
where E.3 follows from the fact that $\hat{m}_{n}(x)\geq Y_{min}\delta\text{ almost surely for all }x\in\mathbb{R}$.
@@
$\leq\mathbb{P}\left[\frac{\sum\limits_{i=1}^{n}K_{1,h}\left(X_{1}-X_{i}\right)\left(Y_{i}-Y_{min}\delta)\right)}{\sum\limits_{i=1}^{n}K_{h}\left(X_{1}-X_{i}\right)}\geq t|X_{1}=x,Y_{1}=y\right]$ (Using E.3)
$=\mathbb{P}\left[\sum\limits_{i=2}^{n}\left\{K_{1,h}\left(x-X_{i}\right)\left(Y_{i}-Y_{min}\delta\right)-tK_{h}\left(x-X_{i}\right)\right\}\geq t\right]$
@@
$\leq\mathbb{P}\left[\sum\limits_{i=2}^{n}\left(Z_{i}-\mathbb{E}\left(Z_{i}\right)\right)\geq t+2h\delta p_{min}(\gamma_{1}+\epsilon)(n-1)\right],$ (E.4 )
@@
$Z_{i}=K_{1,h}\left(x-X_{i}\right)\left(Y_{i}-Y_{min}\delta\right)-(m_{1}(X_{i})+\epsilon)K_{h}\left(x-X_{i}\right).$ (8.3)
@@
Now, we would like to apply Hoeffding’s inequality (See Definition 1) for the random variables $Z_{i}-\mathbb{E}Z_{i},i=2,\ldots n$, therefore consider the following:
@@
    We have $Z_{min}\leq Z_{i}\leq Z_{max}$, almost surely, where $Z_{min}$ and $Z_{max}$ are defined as:
    $Z_{min}=(X_{min}-X_{max})h^{-2}\tau_{1}Y_{min}\delta(1-\delta)-(M_{1}+\epsilon),$
    $Z_{max}=(X_{max}-X_{min})h^{-2}\tau_{2}(Y_{max}-Y_{min}\delta)-(\gamma_{1}+\epsilon)\delta.$
    This follows from Assumptions (A.3), (A.5), (B.1), (B.3), and (B.5). Consequently, assuming $(M_{1}+\epsilon)h^{2}\approx 0$ and $(\gamma_{1}+\epsilon)\delta h^{2}\approx 0$, for small enough $h$, we have:
    $Z_{max}-Z_{min}\approx h^{-2}Q,$ (8.4)
    where $Q=(X_{max}-X_{min})\left[\tau_{2}(Y_{max}-Y_{min}\delta)+\tau_{1}Y_{min}\delta(1-\delta)\right]$.
Therefore, using 1 and 2 above, and applying Hoeffding’s inequality for $u=t+2h\delta p_{min}(\gamma_{1}+\epsilon)(n-1)$, we have
$\mathbb{P}\left[\hat{m}_{1,n}(X_{1})\geq t|X_{1}=x,Y_{1}=y\right]$ $\leq\exp{\frac{-2\left(t+2h\delta p_{min}(\gamma_{1}+\epsilon)(n-1)\right)^{2}}{\left(n-1\right)\left(Z_{max}-Z_{min}\right)^{2}}}$
$\approx\exp{\frac{-2n^{2}h^{2}\left(\frac{t}{nh}+2\delta p_{min}(\gamma_{1}+\epsilon)\frac{(n-1)}{n}\right)^{2}}{(n-1)\left(h^{-2}Q\right)^{2}}}$ (By Equation (8.4))
$\asymp\exp{\frac{-2nh^{2}(2\delta p_{min}(\gamma_{1}+\epsilon))^{2}}{h^{-4}Q^{2}}}$
$=\exp{(-nh^{6}C_{\epsilon})},$ (E.5)
where $C_{\epsilon}=\displaystyle{8\delta^{2}p_{min}^{2}(\gamma_{1}+\epsilon)^{2}Q^{-2},\text{ provided }\gamma_{1}\neq-\epsilon}$.
Now, using Equation E.5 in Equation E.2 we have
$\mathbb{P}\left(\exists X_{i}\in\mathcal{X}_{n}^{t}|X_{i}\notin D_{t-\epsilon}^{+}(h)\right)$ $\leq n\int_{(D_{t-\epsilon}^{+}(h))^{c}}\int_{\mathbb{R}}f(x,y)\exp{(-nh^{6}C_{\epsilon})}dydx$
$=n\exp{(-nh^{6}C_{\epsilon})}\int_{\mathbb{R}}\int_{\mathbb{R}}f(x,y)dydx$
$=ne^{-nh^{6}C_{\epsilon}}$
@@
$Z_{i}^{*}=\left(m_{1}(X_{i})-\epsilon\right)K_{h}(x-X_{i})-K_{1,h}(x-X_{i})\left(Y_{i}-Y_{min}\delta\right).$
@@
$\mathbb{P}\left[\hat{m}_{1,n}(X_{1})\leq t|X_{1}=x,Y_{1}=y\right]\leq\mathbb{P}\left[\sum\limits_{i=2}^{n}\left(Z_{i}^{*}-\mathbb{E}\left(Z_{i}^{*}\right)\right)\geq 2h\delta p_{min}(\gamma_{1}+\epsilon)(n-1)-t\right].$
@@
This completes the proof of Lemma 1. $\Box$
Lemma 2
Under the assumptions (A.1)-(A.5) and (B.1)-(B.5), for every $t>0$, and $\epsilon\in(0,t)$ if $h\equiv h(n)\xrightarrow{}0\text{ as }n\xrightarrow{}\infty\text{ such that }nh^{6}\xrightarrow{}\infty$, then for large enough $n$, we have
$\mathbb{P}(D^{-}_{t+\epsilon}(2h)\subset\hat{D}_{t}(n,h)\subset D^{+}_{t-\epsilon}(2h))\geq 1-3ne^{-C_{\epsilon}nh^{6}}.$ (8.5)
@@
Now, since $\hat{D}_{t}(n,h)\supset\mathcal{X}_{n}^{t}$, using the Lemma 1 we have
@@
$\leq\mathbb{P}\left(\mathcal{X}_{n}^{t}\cap\left(D^{+}_{t-\epsilon}(h)\right)^{c}\neq\varnothing\right)\leq ne^{-nh^{6}C_{\epsilon}}.$ (E.6)
@@
where $F_{X}$ is the distribution function of the regressors. Then, for $0<h<1$, we have
$F_{X}\left(\mathbb{B}_{(1-\gamma)h}(u)\right)=\int\limits_{u-(1-\gamma)h}^{u+(1-\gamma)h}p(x)dx\geq 2(1-\gamma)hp_{min}\geq 2(1-\gamma)h^{6}p_{min}$ (8.8)
@@
$\mathbb{P}\left(\exists u\in\mathcal{I}:\mathbb{B}_{(1-\gamma)h}(u)\cap\mathcal{X}^{t}_{n}=\varnothing\right)\leq N_{\gamma}2\gamma he^{-n2(1-\gamma)h^{6}p_{min}}=N_{\gamma}2\gamma he^{-nh^{6}C_{\gamma}},$ (8.9)
where $C_{\gamma}=2(1-\gamma)p_{min}$. Now, note that the second term in Equation E.8 can be bounded by Lemma 1. Therefore, using the Equation 8.2, we have the following bound for the second term of Equation E.8,
$\mathbb{P}\left(D^{-}_{t+\epsilon}(h)\cap\mathcal{X}_{n}\not\subset\mathcal{X}^{t}_{n}\right)\leq ne^{-C_{\epsilon}nh^{6}}$ (8.10)
Now, using Equation 8.9 and Equation 8.10 in Equation E.8, we have
$\mathbb{P}\left(E_{1}^{c}\right)\leq N_{\gamma}2\gamma he^{-nh^{2}C_{\gamma}}+ne^{-C_{\epsilon}nh^{2}}\leq 2ne^{-C_{\epsilon}nh^{6}},$ (8.11)
@@
Thus, using Equation 8.11 and Equation E.6 in Equation 8.6, we get the desired inequality as in Equation 8.5. This completes the proof of Lemma 2. $\Box$
@@
Therefore, it is enough to prove that $\delta_{B}\left(\widehat{PH}_{k}^{\epsilon}(m_{1}),PH_{k}^{\epsilon}(m_{1})\right)\leq 3\epsilon$ holds with high probability. Let E be the event that the following sequence of inclusions holds for $i=1,\ldots,M_{\epsilon}$:
$D_{t_{i}+\epsilon}\hookrightarrow\hat{D}_{t_{i}}(n,h)\hookrightarrow D_{t_{(i+1)}+\epsilon},$
where $M_{\epsilon}=\min\{i\in\mathbb{N}:t_{i}\geq\gamma_{1}\}$. Recall that $\gamma_{1}=\inf m_{1}(x)$, defined in Assumption (B.5). Note that for all $i\geq M_{\epsilon}$, the level sets $D_{t_{i}}$ remain the same. Now, applying Lemma 2 $M_{\epsilon}$ times for $\hat{D}_{t_{i}}(n,h)$, we have
$\mathbb{P}(E)\geq 1-3nM_{\epsilon}e^{-C_{\epsilon/2}nh^{6}}$
@@
$\mathbb{P}\left(\delta_{B}\left(\widehat{PH}_{k}^{\epsilon}\left(m_{1}\right),PH_{k}\left(m_{1}\right)\right)\leq 5\epsilon\right)\geq 1-3nM_{\epsilon}e^{-C_{\epsilon/2}nh^{6}}.$